OpenStudy (anonymous):

Choose the equation of the line passing through the point (-2, 6) and parallel to y = one halfx - 8. y = one halfx + 7 y = one halfx + 5 y = 2x + 7 y = 2x + 5

OpenStudy (anonymous):

@Princezz @Easyaspi314 @Embryo

OpenStudy (anonymous):

Recall what it means for two lines to be parallel, that means that they both have the same slope, and also recall what the general form of a line is,\[y=mx+b\]where m is the slope, and b is the y-intercept of your graph, now here's where the slope comes into play, in order to find a line that goes through specific points, you plug those points into what's called the point-slope form of the line which is as follows,\[y-y_1=m(x-x_1)\]where your X1 and Y2 are the coordinates of the point you want the line to pass through, and m is the slope, so this is where it's important to know if it's asking for a PARALLEL line or a PERPENDICULAR line, since you want a parallel line, you just plug in the slope you are given, but if it asked for a perpendicular line, then you'd use\[-(m^{-1})\]as your slope instead. Hope that helps.

OpenStudy (anonymous):

im pretty confused not because of your explanation (Which was perfect) But I just don't know what to do

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Alrighty, for the steps, you want to find a line that goes through the points\[(x_1,y_1)=(-2,6)\]and from your original line, you have that the slope of the line is\[m=\frac{ 1 }{ 2 }\]now to find the equation of a line that is parallel, both of these lines will have the same slope, so you've got your points, and now you have your slope, point-slope form tells us that the equation of a line that goes through a point can be written as\[y-y_1=m(x-x_1)\]once you plug in the values for X1,Y1 and m, you solve for y and put it in the form\[y=mx+b\]so in conclusion, you're solving the equation\[y-(6)=\frac{ 1 }{ 2 }(x-(-2))\]for y, and that will give you your parallel line.

OpenStudy (anonymous):

So y = x/2 + 7?

OpenStudy (anonymous):

That's A right?

OpenStudy (anonymous):

Correct !

OpenStudy (anonymous):

Thank You so much!